OpenStudy (anonymous):

A triangle is graphed in the coordinate plane. The vertices of the triangle have coordinates (-3, 1), (1, 1), and (1, -2). What is the perimeter of the triangle? A. 6 B. 7 C. 12 D. 14

OpenStudy (anonymous):

@Jesstho.-.

OpenStudy (anonymous):

You can see two lengths of the triangle, and use Pythagorean Theorem to find the other length. Then, add all sides for perimeter.

OpenStudy (anonymous):

I think B

OpenStudy (anonymous):

I was thinking 7 too. Thanks! :D

OpenStudy (anonymous):

B is incorrect.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Look at the picture. 1 side is 4 units, and one side is 3 units. The triangle in the picture is a right one, so, use Pythagorean Theorem to find the other length(hypotenuse). Ask me if you need more help with Pythagorean Theorem.

OpenStudy (anonymous):

Can you walk me through it? I can't remember how to do it.

OpenStudy (anonymous):

Pythagorean Theorem: \[a ^{2}+b ^{2}=c ^{2}\]\[a ^{2}=9\]\[b ^{2}=16\]\[16+9=25\]\[\sqrt{25}=5\]So: a=3 b=4 and c, the missing value=5 Now, add all the sides. 3+4+5=12

OpenStudy (anonymous):

Oh okay! Thank you so so SO much!!

OpenStudy (anonymous):

Your welcome.